Write the equation of the line that passes through the points (2,7) and (8,-1).

Answer by nyc_function(2741)

Let m = slope
m = (-1 - 7)/8 - 2)
m = -8/6
m = -4/3
We now use the point-slope formula to find the equation.
The formula is y - y1 = m(x - x1).
We plug the slope and one of the points given to you.
I will use (2,7) but you can use the other point if your want to.
y - 7 = (-4/3)(x - 2)
y - 7 = (-4/3)(x) + 8/3
y = (-4/3)(x) + 8/3 + 7
y = (-4/3)(x) + 29/3
Done!